OpenStudy (ddcamp):

Find the Maclaurin series representation of 2*e^(2x+1) and use it to approximate 2*e^5 to 6 decimal places.

OpenStudy (ddcamp):

I know that the Maclaurin series for e^x is: \[\sum_{n=0}^{\infty} \frac{ x^{n} }{ n! }\] So would this series be: \[2\sum_{n=0}^{\infty} \frac{ (2x+1)^{n} }{ n! }?\]

OpenStudy (anonymous):

yeah i guess you could expand that way or else just expand

OpenStudy (ddcamp):

OK. Though honestly I don't think I need to do anything other than just plug 5 into the original series. XD silly calculus problems.